A board, 39 cm long is cut into three pieces such that the second piece is twice as long as the first and the third is 4 cm longer than the second. Find the length of the shorter piece.
Answer by addingup(3677)

3 pieces: x, y, z
the second, y, is 2x or y = 2x
the 3rd, z, is 4cm longer than y, or z = y+4
x+y+z = 39 substituting:
x+2x+y+4 = 39 OK, this won't work because x and y don't mix. But since we know that y = 2x, we can add 4 to this y-equivalent, like this:
x+2x+2x+4 = 39
Now we've got it. Add up the x:
5x+4 = 39
5x = 35
x = 7
-------------------------------------
Check:
x = 7
y = 2x = 14
z = 2x+4 = 18
7+14+18 = 39 Correct